OpenStudy (anonymous):

Give exact and approximate solutions to three decimal places. (x+7)^2=36

myininaya (myininaya):

take square root of both sides then subtract 7 on both sides

OpenStudy (anonymous):

and don't forget to write both \[x+7=6\] or \[x+7=-6\]

OpenStudy (amistre64):

id +- a sqrt(36) then subtract out a 7

OpenStudy (anonymous):

that's easy for you to say

OpenStudy (amistre64):

\[(\pm \sqrt{36}-7+7)^2 = 36\]